OpenStudy (anonymous):

multiply x²+ 2x-15 7x-14 --------- * ------ x²+4x-5 x²-4 write answer in simplest form.

OpenStudy (anonymous):

factor everything first so you can get rid of terms can you do this?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[=((x ^{2}+2x-15)(7x-14))/((x ^{2}+4x-5)(x ^{2}-4))\] \[=((x+5)(x-3)\times7(x-2))/((x+5)(x-1)\times(x-2)(x+2))\] when\[ x \neq -5\] \[x \neq 2\] \[=7(x-3)/(x ^{2}+x-2)\]